OpenStudy (anonymous):

Determine whether each sequence could be arithmetic. If so give the common difference. 17, 34, 51, 68,

jimthompson5910 (jim_thompson5910):

Take each term and subtract off the previous term 34 - 17 = ??? 51 - 34 = ??? 68 - 51 = ???

jimthompson5910 (jim_thompson5910):

if each difference is the same, then you have an arithmetic sequence and this difference is the common difference

OpenStudy (anonymous):

34-17=17 51-34=17 68-51=17

jimthompson5910 (jim_thompson5910):

so that shows you this sequence is arithmetic

jimthompson5910 (jim_thompson5910):

so the common difference is 17

OpenStudy (anonymous):

Thanks(:

jimthompson5910 (jim_thompson5910):

np